# Learning on LLM Output Signatures for gray-box LLM Behavior Analysis

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse applications, yet their internal mechanisms remain poorly understood. This gap in understanding is particularly relevant in critical tasks like Hallucination Detection (HD) [45, 28, 17, 20, 40] and Data Contamination Detection (DCD) [6, 41, 56]: in these contexts, determining whether an LLM is fabricating information or has been exposed to specific training data is crucial for deployment safety and reliability.
@@
Figure 1: A visualization of the LLM Output Signature (LOS). Left: The LLM processes the input “What does the cat chase?” and generates the output “A big mouse”. Right: The corresponding query/response Token Distribution Sequence (TDS) – full next-token distributions over the sequence – and Actual Token Probabilities (ATP), which represent the probabilities of tokens in the actual sequence (for the query) and the sampled token probabilities (for the response). We propose learning directly from this unified LOS representation to analyze LLM behavior.
@@
Previous work on LLM analysis has relied heavily on probing techniques that require restrictive white-box access to model internals [3, 35, 16, 15, 39]. Gray-box methods relax these assumptions by operating only on LLM outputs. Existing gray-box approaches typically analyze just the sequence of probabilities assigned to tokens that appear in the relevant input or output token sequence – a vector we term Actual Token Probabilities (ATP) [14, 22, 48, 18]. However, these methods, often based on heuristics, overlook the information contained in the complete Token Distribution Sequence (TDS) – a matrix holding the next-token probability distribution at each generation step, see Figure 1. This limitation can mask crucial differences in model behavior even at the level of a single time step. E.g., consider a model generating a token with probability 0.5 in two scenarios: in one case, the remaining next-token probability mass is concentrated on a single alternative $(0.5,0.5,0,...,0)$, while in the other it is spread across many tokens: $(0.5,0.01,...,0.01)$. These distributions suggest very different levels of model uncertainty, yet ATP-based approaches would treat them identically. Similarly, an ATP value of 0.1 at a certain time step could indicate either high uncertainty (if it is the highest probability in a diffused distribution) or strong evidence against the token (if it is a low-ranking probability in a peaked distribution). A recent promising approach [56] used some TDS information through heuristics, but a principled framework to utilize this data lacks.
@@
Our approach. We argue that a successful gray-box approach should leverage both ATP and TDS, together forming what we term the LLM Output Signature (LOS) (Figure 1) – the complete observable representation of LLM behavior in the gray-box setup. Instead of relying on heuristics, we treat LOS as a sequential, high-dimensional and structured data modality on which we apply principled deep learning techniques. We propose LOS-Net, a lightweight transformer encoder111 Less than 1M parameters. that operates on an effective encoding of ATP, TDS, and their interactions. From a theoretical perspective, we show that LOS-Net can provably approximate a broad class of functions applied to the LOS of any LLM, subsuming many recent approaches [14, 22, 48, 18, 41, 56]. Through a comprehensive empirical study on the tasks of DCD and HD, we demonstrate that the information gap between using the complete LOS and relying solely on ATP is substantial. We improve over all considered baselines across both tasks – often by a significant margin. Notably, LOS-Net exhibits promising dataset-level transfer and strong cross-model generalization, suggesting it can capture universal patterns in LLM behavior. Importantly, the cross-model transfer abilities of LOS-Net suggest its viable application to impactful real-world tasks such as copyright-infringement detection over closed-source LLMs, as indicated by our results on the BookMIA benchmark [41].
@@
Contributions are summarized as follows: (1) we introduce LOS as a suitable representation for analyzing LLM behavior, (2) we develop an effective learning framework for the LOS data modality, (3) we show this unifies and generalizes previous approaches, (4) we demonstrate it achieves superior performance across models, datasets, and tasks, and (5) exhibits strong empirical evidence for cross-model generalization and promising cross-dataset transfer abilities. The proposed LOS-Net proves effective for both HD and DCD, and its flexibility suggests broader potential for similar tasks while paving the way for foundational approaches to modeling LLM behaviors.
@@ 2 Related Work
We review background and related work on DCD and HD, focusing on studies leveraging logits or output probabilities. Given the breadth of research, we highlight the most relevant works for our setup and refer interested readers to Appendix D for further details on these tasks.
@@
Hallucination Detection. Hallucination detection has been studied as a means of enabling selective intervention, allowing LLMs to prevent fabricated outputs only when necessary [44, 54, 47]. Recently, [35] showed that training a classifier on top of LLMs’ hidden states (intermediate activations) is highly effective for hallucination detection. However, this method operates under the white-box assumption, requiring full access to the model’s internal components. In contrast, our paper explores a more constrained (gray-box) setting.
@@ 3.1 Notation and Problem Formulation
Figure 2: LLM processing pipeline. A token sequence $\vec{s}$ is processed by an LLM $f$ and generates full TDSs $\textbf{X}_{s},\textbf{X}_{g}$ corresponding to the input $\vec{s}$ and the response $\vec{g}$.
Let $f$ denote a pretrained LLM, and $\vec{s}$ refer to a text input to $f$ consisting of $n$ tokens. When queried with $\vec{s}$, the LLM $f$ produces outputs $\mathbf{X}_{s}=f(\vec{s})$, i.e., a matrix in $\mathbb{R}^{n\times V}$ of next-token probabilities for each token in $\vec{s}$, where $V$ is the size of the token vocabulary. We define the LLM response to be $\vec{g}$ consisting of $m$ tokens generated using $f$’s outputs in $\mathbf{X}_{g}\in\mathbb{R}^{m\times V}$ (and $\mathbf{X}_{s}$). We refer to $\mathbf{X}_{s}$ or $\mathbf{X}_{g}$ as Token Distribution Sequences (TDS). See Figure 2.
We also define $\mathbf{p}_{s}\in\mathbb{R}^{n},\mathbf{p}_{g}\in\mathbb{R}^{m}$, which holds the probabilities associated with the actual tokens appearing in $\vec{s},\vec{g}$ respectively. We denote these as the Actual Token Probabilities (ATP). Specifically, $(\mathbf{p}_{s})_{i}\coloneqq\mathbf{X}_{i,v}$ where $v\in\{1,\dots,V\}$ is the token used in the $i+1$ place in the sequence $\vec{s}$ and similarly for $\vec{g}$. See Figure 1 for an illustration.
We call the pairs $(\mathbf{X}_{s},\mathbf{p}_{s})$ or $(\mathbf{X}_{g},\mathbf{p}_{g})$ the LLM Output Signature (LOS). For DCD, we analyze input sequences using $(\mathbf{X}_{s},\mathbf{p}_{s})$ since our interest lies in how the model processes the input text $\vec{s}$. In contrast, for HD, we use $(\mathbf{X}_{g},\mathbf{p}_{g})$ as we need to analyze the model’s generated response. We will sometimes use $(\mathbf{X},\mathbf{p})$ if the distinction between the tasks is irrelevant, and use $N$ as the sequence length.
Problem Statement. LOS elements, along with their associated annotations depending on the task of interest, can be gathered into datasets $D=\{\big{(}(\mathbf{X},\mathbf{p})_{i},y_{i}\big{)}\}_{i=1}^{\ell}$
where supervised learning problems can be instantiated. Our goal in this paper is to propose a neural architecture that can effectively utilize the complete LOS to solve tasks such as DCD, HD (see inset), or any other classification problem defined thereon.
@@ 3.2 Our Approach
Our approach consists of three main steps. Given an input $(\mathbf{X},\mathbf{p})$: (1) The probability distributions in TDS ($\mathbf{X}$) are sorted independently and sliced to only include the top $K$ ones at each time step, obtaining $\mathbf{X}^{\prime}$ ; (2) A learnable Rank Encoding $\text{RE}(\mathbf{X},\mathbf{p})$ is concatenated to $\mathbf{X}^{\prime}$ to capture relative probability information;
(3) The resulting representation is processed by a lightweight transformer architecture, yielding the desired output. See inset illustration. In the remainder of this section, we provide a detailed explanation of each component.
@@
Utilizing $\mathbf{X}$ may pose significant challenges due to three key factors. (1) Complexity: The vocabulary tensor can be extremely large in real-world scenarios. For instance, Liang et al. [27] (XLM-V) reported a vocabulary size of 1M tokens, which, for a small batch of documents and popular context sizes, would already entail processing a tensor of tens (or hundreds) of GBs. (2) Transferability: Vocabulary size and order may significantly vary between LLMs, something which can complicate transfer learning – e.g., training on one LLM and testing on another with a different vocabulary size; (3) Limited Access: In certain LLMs, such as those released by OpenAI, the output tensor $\mathbf{X}$ is only partially accessible, with APIs only exposing the (log-) probabilities for a small number of most likely tokens.
To tackle these challenges, we propose selecting, for each row of $\mathbf{X}$, a fixed number of elements. Specifically, we preprocess $\mathbf{X}$ by sorting each row independently and selecting the top $K$ probabilities, as follows:
@@
The tensor $\mathbf{X}^{\prime}$ provides a comprehensive description of the LLM’s output, but does not encode an important source of information: the probability $\mathbf{p}$ of the actual tokens appearing in the sequence, i.e, the ATP. The importance of this feature both in DCD and HD has already been demonstrated by a large body of prior work that operated only on this information [41, 14, 22, 48, 18]. Taking inspiration from these, we do also include ATPs as inputs to our architecture. However, we further complement these probabilities with additional information which allows us to contextualize them with respect to the whole TDS, i.e., $\mathbf{X}$. Specifically, we argue that valuable information is encoded in the rank (position) of the ATP within the vocabulary-wide (sorted) sequence of token probabilities. This information reveals both the model’s generation patterns and potential mismatches between predicted and actual tokens. The rank of the $i$-th token in the sequence is defined as: $r_{i}(\mathbf{X},\mathbf{p})=\sum_{v=1}^{V}\mathbb{I}(\mathbf{X}_{i,v}>p_{i}),$ where $\mathbb{I}(\cdot)$ is the indicator function.
We encode the rank in a way to make this feature more amenable for learning, while still maintaining enough expressivity. Specifically, we first scale the rank between $[-1,1]$, obtaining $\mathbf{r}^{\text{scaled}}$. Then, we construct the following learnable rank encoding [^2] For certain DC datasets, we used a lookup table for Rank encoding, where the index corresponds to $r_{i}$ and the value is an embedding.,
$\text{RE}(\mathbf{X},\mathbf{p})=\mathbf{p}\otimes\left(\mathbf{r}^{\text{scaled}}\cdot\mathbf{w}_{1}+\mathbf{w}_{2}\right),$ (2)
where $\otimes$ is an outer product, and $\mathbf{w}_{1},\mathbf{w}_{2}$ are learnable parameters in $\mathbb{R}^{d}$. As a result, $\text{RE}(\mathbf{X},\mathbf{p})$ is in $\mathbb{R}^{N\times d}$. Importantly, the multiplication by $\mathbf{p}$ makes sure that the rank encoding and the TDS are in similar scales, especially when using log probabilities or logits.
@@
Given the preprocessed TDS $\mathbf{X}^{\prime}$, and the developed rank encodings $\text{RE}(\mathbf{X},\mathbf{p})$, our approach applies an encoder-only transformer model $\mathcal{T}$ with learnable positional encodings in the temporal dimension [49], on the token-wise linear projection of $\mathbf{X}^{\prime}$, concatenated to the rank encodings $\text{RE}(\mathbf{X}^{\prime},\mathbf{p})$,
@@ 4 Generalization of Previous Approaches
In this section, we demonstrate that LOS-Net generalizes several leading existing methods through specific weight configurations. This ensures that our architecture can theoretically match these methods, while in practice significantly outperforming them, as demonstrated in our experiments. As already mentioned, prior research has introduced various methods for analyzing LLMs based on their output probabilities [14, 22, 48, 18], with many approaches focusing on the ATPs. We note that many of these methods assume the form of statistics calculated over the whole sequence processed by the LLM. Recent, more sophisticated approaches aggregate these probabilities only for some of the tokens in the sequence, dynamically chosen based on features computed on the set of ATPs [41, 56], as we illustrate below.
@@
This highlights a general pattern: that of computing a global score by aggregating token-wise values meeting a (dynamic) “acceptance” condition, a form of ‘gating’. To unify the aforementioned baselines under a common framework, we formalize this pattern in a family of functions, defined next.
Gated Scoring Functions (GSFs). We define the family of ‘Gated Scoring Functions’ (GSF) as the set of functions that score LOSs by aggregating token-wise scores across the input sequence whenever their confidence values exceed a (possibly adaptive) threshold. GSFs are described in terms of the following three components: (1) A confidence function $\kappa:\mathbb{R}^{N\times k}\times\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ that assigns confidence values to each token in the sequence; (2) A threshold function $T:\mathbb{R}^{N\times k}\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ that determines an acceptance criterion; and (3) A weight function $g:\mathbb{R}^{N\times k}\times\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ that assigns importance scores to tokens. Given a LOS $(\mathbf{X},\mathbf{p})$, a GSF computes a global score $R(\mathbf{X},\mathbf{p})$ as follows:
$F(\mathbf{X},\mathbf{p})_{i}$ =$=\begin{cases}g(\mathbf{X}^{\prime},\mathbf{p})_{i},&\text{if }\kappa(\mathbf{X}^{\prime},\mathbf{p})_{i}\geq T(\mathbf{X}^{\prime},\mathbf{p}),\\ 0,&\text{otherwise},\end{cases}$ (5)
$R(\mathbf{X},\mathbf{p})$ $=\sum_{i=1}^{N}F(\mathbf{X},\mathbf{p})_{i},$ (6)
@@
It is easy to see, e.g., how MinKis implemented as a GSF. For a sequence length of $N$, it suffices to choose:
$T(\mathbf{X}^{\prime},\mathbf{p})$ $=-\text{perc}(\mathbf{p},K)=-\big{(}\text{sort}(\mathbf{p})_{\left\lceil\frac{K}{100}\cdot N\right\rceil}\big{)},$
$\kappa(\mathbf{X}^{\prime},\mathbf{p})$ $=-\mathbf{p},\quad g(\mathbf{X}^{\prime},\mathbf{p})=\frac{\log\mathbf{p}}{\left\lceil\frac{K}{100}\cdot N\right\rceil}.$
We refer readers to Appendix B for a proof of Proposition 4.1 and more details on how other baselines are implemented.
@@
Proposition 4.2 (Our model can approximate Equation 6).
Assume a maximal possible vocabulary size $V_{\text{max}}$ and a maximal context size $N_{\text{max}}$. Let $\mathcal{X}\times\mathcal{M}\subseteq\mathbb{R}^{N_{\text{max}}\times V_{\text{max}}}\times\mathbb{R}^{N_{\text{max}}}$ represent a compact subset in the LOS. For any measurable $\kappa:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}^{N_{\text{max}}}$, measurable $T:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}$, measurable and integrable weight function $g:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}^{N_{\text{max}}}$, and for any $\epsilon>0$, there exists a set of parameters $\theta$ such that our model $h_{\theta}:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}$ satisfies $\left\lVert h_{\theta}-R\right\rVert_{L_{1}}<\epsilon$ where $\left\lVert\cdot\right\rVert_{L_{1}}$ denotes the $L_{1}$ norm.
The complete proof is given in Appendix B. To prove this result, we build on existing universality results for approximating continuous functions with Transformers [55], by showing that our (generally non-continuous) target functions can be approximated by continuous functions. The implications of this proposition are interesting: as long as the LOS space of interest lies within a compact domain – an assumption inherently satisfied when using, e.g., probabilities [^3] For logits or log probabilities, clamping ensures the compactness assumption. – our model can approximate the general GSF given in Equation 6 of LOSs of any LLM under mild conditions on $\kappa$, $T$, and $g$, potentially generalizing across LLMs as our approximation result considers a predefined LOS domain. In Section 5 we show that our trained models can indeed be applied successfully out-of-the-box on LOSs from different LLMs. We note that Proposition 4.2 cannot be generally extended to $L_{\infty}$ due to the discontinuity of GSFs. The practical relevance of Proposition 4.2, is underscored by the following corollary:
@@
The above states that well-established, successful baselines from the literature (see class $\mathcal{B}$ in Proposition 4.1) can be approximated by LOS-Net. The proof for Corollary 4.3 follows from Propositions 4.1 and 4.2, see Appendix B for the complete proof.

### v2

# Learning on LLM Output Signatures for Gray-Box Behavior Analysis

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse applications, yet their internal mechanisms remain poorly understood. This gap in understanding is particularly relevant in critical tasks like Hallucination Detection (HD) [45, 28, 17, 20, 40] and Data Contamination Detection (DCD) [6, 41, 58], where determining whether an LLM is fabricating information or has been exposed to specific training data is crucial for safe and reliable deployment.
@@
Figure 1: Left: The LLM processes the input “What does the cat chase?” and generates the output “A big mouse”. Right: The corresponding query/response (“resp.”) Token Distribution Sequences (TDS) and Actual Token Probabilities (ATP), together constituting the LLM Output Signature (LOS). We propose to analyze LLM behavior by learning directly from this unified representation.
@@
Previous work on LLM analysis has relied heavily on probing techniques that require restrictive white-box access to model internals [3, 35, 16, 15, 39]. Gray-box methods relax these assumptions by operating only on LLM outputs. Existing gray-box approaches typically analyze just the sequence of probabilities assigned to tokens that appear in the relevant input or output token sequence – a vector we term Actual Token Probabilities (ATP) [14, 22, 49, 18]. However, these methods, often based on heuristics, overlook the information contained in the complete Token Distribution Sequence (TDS) – a matrix holding the next-token probability distribution at each generation step, see Figure 1. This limitation can mask crucial differences in model behavior even at the level of a single time step. E.g., consider a model generating a token with probability 0.5 in two scenarios: in one case, the remaining next-token probability mass is concentrated on a single alternative $(0.5,0.5,0,...,0)$, while in the other it is spread across many tokens: $(0.5,0.01,...,0.01)$. These distributions suggest very different levels of model uncertainty, yet ATP-based approaches would treat them identically. Similarly, an ATP value of 0.1 at a certain time step could indicate either high uncertainty (if it is the highest probability in a diffused distribution) or strong evidence against the token (if it is a low-ranking probability in a peaked distribution). A recent promising approach [58] used some TDS information using heuristics, but a principled framework to utilize this data is still lacking.
@@
Our approach. We argue that a successful gray-box approach should leverage both ATP and TDS, together forming what we term the LLM Output Signature (LOS) (Figure 1) – the complete observable representation of LLM behavior in the gray-box setup. Instead of relying on heuristics, we treat LOS as a sequential, high-dimensional and structured data modality on which we apply principled deep learning techniques. We propose LOS-Net, a lightweight transformer encoder [^1] Around 1M parameters. operating on an effective encoding of ATP, TDS, and their interactions. We prove that LOS-Net can approximate a broad class of functions applied to the LOS of any LLM, subsuming many recent approaches [14, 22, 49, 18, 41, 58]. Our comprehensive empirical study on DCD and HD demonstrates a substantial information gap between using the complete LOS and relying solely on ATP. Notably, LOS-Net improves over all considered baselines across both tasks, often by a significant margin. Crucially, our architecture is extremely efficient, with detection times of $\approx 10^{-5}$s per instance, making it a compelling approach for applications such as on-line error detection for guided-generation. LOS-Net also exhibits promising dataset-level transfer and strong cross-model generalization, the latter suggesting its viable application to real-world tasks such as copyright-infringement detection over closed-source LLMs (see, e.g., our results on the BookMIA benchmark [41] in Section 5.2).
@@
Contributions are summarized as follows: (1) we introduce LOS as a suitable representation for analyzing LLM behavior, (2) we develop an effective and a learning framework for the LOS data modality, (3) we show this unifies and generalizes previous approaches, (4) we demonstrate it achieves superior performance across models, datasets, and tasks, and (5) exhibits strong empirical evidence for cross-model generalization and promising cross-dataset transfer abilities. The proposed LOS-Net proves effective for both HD and DCD, and its flexibility suggests broader potential for similar tasks while paving the way for foundational approaches to modeling LLM behaviors.
@@ 2 Related Work
We review background and related work on DCD and HD, focusing on studies leveraging logits or output probabilities. Given the breadth of research, we highlight the most relevant works for our setup and refer interested readers to Appendix C for further details on these tasks.
@@
Hallucination Detection has been studied as a means of enabling selective intervention, allowing LLMs to prevent fabricated outputs only when necessary [44, 56, 48]. Recently, [35] showed that training a classifier on top of LLMs’ hidden states is highly effective for hallucination detection. However, this method operates under the white-box assumption, requiring full access to the model’s internal components. In contrast, our paper explores a more constrained (gray-box) setting.
@@ 3.1 Notation and Problem Formulation
Figure 2: LLM processing pipeline. Token sequence $\vec{s}$ is processed by an LLM $f$, generating full TDSs $\textbf{X}_{s},\textbf{X}_{g}$ for input $\vec{s}$ and response $\vec{g}$.
Let $f$ denote a pretrained LLM, and $\vec{s}$ refer to a text input to $f$ consisting of $n$ tokens. When queried with $\vec{s}$, the LLM $f$ produces outputs $\mathbf{X}_{s}=f(\vec{s})$, i.e., a matrix in $\mathbb{R}^{n\times V}$ of next-token probabilities for each token in $\vec{s}$, where $V$ is the size of the token vocabulary. We define the LLM response to be $\vec{g}$ consisting of $m$ tokens generated using $f$’s outputs in $\mathbf{X}_{g}\in\mathbb{R}^{m\times V}$ (and $\mathbf{X}_{s}$). We refer to $\mathbf{X}_{s}$ or $\mathbf{X}_{g}$ as Token Distribution Sequences (TDS). See Figure 2. We also define $\mathbf{p}_{s}\in\mathbb{R}^{n},\mathbf{p}_{g}\in\mathbb{R}^{m}$, which holds the probabilities associated with the actual tokens appearing in $\vec{s},\vec{g}$ respectively. We denote these as the Actual Token Probabilities (ATP). Specifically, $(\mathbf{p}_{s})_{i}\coloneqq\mathbf{X}_{i,v}$ where $v\in\{1,\dots,V\}$ is the token used in the $i+1$ place in the sequence $\vec{s}$ and similarly for $\vec{g}$. See Figure 1 for an illustration. We call the pairs $(\mathbf{X}_{s},\mathbf{p}_{s})$ or $(\mathbf{X}_{g},\mathbf{p}_{g})$ the LLM Output Signature (LOS). For DCD, we analyze input sequences using $(\mathbf{X}_{s},\mathbf{p}_{s})$ since our interest lies in how the model processes the input text $\vec{s}$. In contrast, for HD, we use $(\mathbf{X}_{g},\mathbf{p}_{g})$ as we need to analyze the model’s generated response. We may use $(\mathbf{X},\mathbf{p})$ if the distinction between the tasks is irrelevant, and use $N$ as the sequence length.
Problem Statement. LOS elements, along with their associated annotations depending on the task of interest, can be gathered into datasets $D=\{\big{(}(\mathbf{X},\mathbf{p})_{i},y_{i}\big{)}\}_{i=1}^{\ell}$ where supervised learning problems can be instantiated. Our goal in this paper is to propose a neural architecture that can effectively utilize the complete LOS to solve tasks such as DCD, HD, or any other classification problem thereon.
@@ 3.2 Our Approach
Our approach consists of three main steps. Given an input $(\mathbf{X},\mathbf{p})$: (1) The probability distributions in TDS ($\mathbf{X}$) are sorted independently and sliced to only include the top $K$ ones at each time step, obtaining $\mathbf{X}^{\prime}$ ; (2) A learnable Rank Encoding $\text{RE}(\mathbf{X},\mathbf{p})$ is concatenated to $\mathbf{X}^{\prime}$ to capture relative probability information; (3) The resulting representation is processed by a lightweight transformer architecture, yielding the desired output. See illustration in Appendix E, Figure 9. In the remainder of this section, we provide a detailed explanation of each component.
@@
Utilizing $\mathbf{X}$ may pose significant challenges due to three key factors. (1) Complexity: The vocabulary tensor can be extremely large in real-world scenarios. For instance, Liang et al. [27] (XLM-V) reported a vocabulary size of 1M tokens, which, for a small batch of documents and popular context sizes, would already entail processing a tensor of tens (or hundreds) of GBs. (2) Transferability: Vocabulary size and order may significantly vary between LLMs, something which can complicate transfer learning – e.g., training on one LLM and testing on another with a different vocabulary size; (3) Limited Access: In certain LLMs, such as those released by OpenAI, the output tensor $\mathbf{X}$ is only partially accessible, with APIs only exposing the (log-) probabilities for a small number of most likely tokens. To tackle these challenges, we propose selecting, for each row of $\mathbf{X}$, a fixed number of elements. Specifically, we preprocess $\mathbf{X}$ by sorting each row independently and selecting the top $K$ probabilities, as follows:
@@
The tensor $\mathbf{X}^{\prime}$ provides a comprehensive description of the LLM’s output, but does not encode an important source of information: the probability $\mathbf{p}$ of the actual tokens appearing in the sequence, i.e, the ATP. The importance of this feature both in DCD and HD has already been demonstrated by a large body of prior work that operated only on this information [41, 14, 22, 49, 18]. Taking inspiration from these, we do also include ATPs as inputs to our architecture. However, we further complement these probabilities with additional information which allows us to contextualize them with respect to the whole TDS, i.e., $\mathbf{X}$. Specifically, we argue that valuable information is encoded in the rank (position) of the ATP within the vocabulary-wide (sorted) sequence of token probabilities. This information reveals both the model’s generation patterns and potential mismatches between predicted and actual tokens. The rank of the $i$-th token in the sequence is defined as: $r_{i}(\mathbf{X},\mathbf{p})=\sum_{v=1}^{V}\mathbb{I}(\mathbf{X}_{i,v}>p_{i}),$ where $\mathbb{I}(\cdot)$ is the indicator function. We encode the rank in a way to make this feature more amenable for learning, while still maintaining enough expressivity. Specifically, we first scale the rank between $[-1,1]$, obtaining $\mathbf{r}^{\text{scaled}}$. Then, we construct the following learnable rank encoding [^2] For certain DC datasets, we used a lookup table for Rank encoding, where the index corresponds to $r_{i}$ and the value is an embedding.,
$\text{RE}(\mathbf{X},\mathbf{p})=\mathbf{p}\odot\mathbf{r}^{\text{scaled}}\cdot\mathbf{w}_{1}+\mathbf{p}\cdot\mathbf{w}_{2},$ (2)
where $\odot$ is the hadamard product, and $\mathbf{w}_{1},\mathbf{w}_{2}$ are learnable parameters in $\mathbb{R}^{d}$. As a result, $\text{RE}(\mathbf{X},\mathbf{p})$ is in $\mathbb{R}^{N\times d}$. Importantly, the multiplication by $\mathbf{p}$ makes sure that the rank encoding and the TDS are in similar scales, especially when using log probabilities or logits.
@@
Given the preprocessed TDS $\mathbf{X}^{\prime}$ and the developed rank encodings $\text{RE}(\mathbf{X},\mathbf{p})$, we first linearly project $\mathbf{X}^{\prime}$, concatenate it with $\text{RE}(\mathbf{X},\mathbf{p})$, and then feed it to an encoder-only transformer $\mathcal{T}$ with learnable positional encodings, operating in the temporal dimension [50]:
@@ 4 Generalization of Previous Approaches
Here, we demonstrate that LOS-Net generalizes several leading existing methods through specific weight configurations. This ensures that our architecture can theoretically match these methods –while significantly outperforming them in practice, as shown in our experiments. Proofs are enclosed in Appendix A. As already mentioned, prior research has introduced various methods for analyzing LLMs based on their output probabilities [14, 22, 49, 18], with many approaches focusing on the ATPs. We note that many of these methods assume the form of statistics calculated over the whole sequence processed by the LLM. Recent, more sophisticated approaches aggregate these probabilities only for some of the tokens in the sequence, dynamically chosen based on features computed on the set of ATPs [41, 58], as we illustrate below.
@@
This highlights a general pattern: that of computing a global score by aggregating token-wise values meeting a (dynamic) “acceptance” condition, a form of “gating”. To unify the aforementioned baselines under a common framework, we formalize this pattern via a family of functions (see next).
Gated Scoring Functions (GSFs). We define the family of Gated Scoring Functions (GSF) as the set of functions scoring LOSs by aggregating token-wise scores across the input sequence whenever their confidence values exceed a (possibly adaptive) threshold. GSFs are described in terms of the following components: (1) A confidence function $\kappa:\mathbb{R}^{N\times k}\times\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ that assigns confidence values to each token in the sequence; (2) A threshold function $T:\mathbb{R}^{N\times k}\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ that determines an acceptance criterion; and (3) A weight function $g:\mathbb{R}^{N\times k}\times\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ that assigns importance scores to tokens. Given a LOS $(\mathbf{X},\mathbf{p})$, a GSF computes a global score $R(\mathbf{X},\mathbf{p})$ as follows:
$F(\mathbf{X},\mathbf{p})_{i}$ =$=\begin{cases}g(\mathbf{X}^{\prime},\mathbf{p})_{i},&\text{if }\kappa(\mathbf{X}^{\prime},\mathbf{p})_{i}\geq T(\mathbf{X}^{\prime},\mathbf{p}),\\ 0,&\text{otherwise},\end{cases}\qquad R(\mathbf{X},\mathbf{p})$ $=\sum_{i=1}^{N}F(\mathbf{X},\mathbf{p})_{i},$ (5)
@@
It is easy to see, e.g., how MinKis implemented as a GSF. For a sequence length of $N$, it suffices to choose: $T(\mathbf{X}^{\prime},\mathbf{p})=-\text{perc}(\mathbf{p},K)=-\big{(}\text{sort}(\mathbf{p})_{\left\lceil\frac{K}{100}\cdot N\right\rceil}\big{)},\quad\kappa(\mathbf{X}^{\prime},\mathbf{p})=-\mathbf{p},\quad g(\mathbf{X}^{\prime},\mathbf{p})=\frac{\log\mathbf{p}}{\left\lceil\frac{K}{100}\cdot N\right\rceil}.$ Refer to Appendix A for more details on how other baselines are implemented.
@@
Proposition 4.2 (LOS-Net can approximate Equation 5).
Assume maximal possible vocabulary size $V_{\text{max}}$ and context size $N_{\text{max}}$. Let $\mathcal{X}\times\mathcal{M}\subseteq\mathbb{R}^{N_{\text{max}}\times V_{\text{max}}}\times\mathbb{R}^{N_{\text{max}}}$ represent a compact subset in the LOS. For any measurable $\kappa:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}^{N_{\text{max}}}$, measurable $T:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}$, measurable and integrable weight function $g:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}^{N_{\text{max}}}$, and for any $\epsilon>0$, there exists a set of parameters $\theta$ such that our model $h_{\theta}:\mathcal{X}\times\mathcal{M}\rightarrow\mathbb{R}$ satisfies $\left\lVert h_{\theta}-R\right\rVert_{L_{1}}<\epsilon$ where $\left\lVert\cdot\right\rVert_{L_{1}}$ denotes the $L_{1}$ norm.
To prove this result, we build on existing universality results on approximating continuous functions with Transformers [57], showing that our (generally non-continuous) target functions can be approximated by continuous functions. Importantly, Proposition 4.2 implies that, as long as the LOS space of interest lies within a compact domain [^3] This is inherently satisfied when using probabilities; or via clamping in the case of logits or log-probs., our model can approximate the general GSF in Equation 5 of LOSs for any LLM under mild conditions on $\kappa$, $T$, and $g$, potentially generalizing across LLMs (as our result considers a predefined LOS domain). In Section 5 we show that our trained models can indeed be applied successfully out-of-the-box on LOSs from different LLMs. Note that Proposition 4.2 cannot be generally extended to $L_{\infty}$ due to the discontinuity of GSFs.
The practical relevance of Proposition 4.2, is underscored by the following:
@@
The above states that well-established, successful baselines from the literature (see class $\mathcal{B}$ in Proposition 4.1) can be approximated by LOS-Net. The proof follows from Propositions 4.1 and 4.2.